OpenStudy (anonymous):

What is the 13th term of this arithmetic sequence? 132, 135, 138, 141, …

OpenStudy (anonymous):

a 168 b 172 c 176 d 179 is it d?

OpenStudy (studygurl14):

135 - 132 = 3 138 - 135 = 3 Add 3 to previous number. They give you 4. What can you do to find the answer?

OpenStudy (anonymous):

add 4?? sorry im confused:/

OpenStudy (studygurl14):

no I mean they give you the first 4 numbers

OpenStudy (anonymous):

make a table 132, 135, 138, 141 .... ? 1 2 3 4 ....13 to get from one elementh to then next you need to add so lets do that 132+x=135 x=3 when solve for x so lets continue our table 132, 135, 138, 141 .... ? 1 2 3 4 ....13 <ith element 0*3 1*3 2*3 3*3 12*3 the 3rd row says we added i*3 where i is current element-1 we can see the pattern emerge that says we add 3rd row to original number in the sequence like so ((13-1)*3)+132 12*3+132 36+132 168

OpenStudy (anonymous):

oh ok ok .. now i get it lol thank you!

OpenStudy (anonymous):

gosh it was easier than what i thought it was lol thanks guys